OpenStudy (anonymous):

ok use continuity to evaluate the limits lim as x approaches 1 (8+sqrtx)/(sqrt 8+x)...

OpenStudy (anonymous):

now i got the asnwer to be 8 +\[\sqrt{1}\]/ 3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

just plug in 1

OpenStudy (anonymous):

oh i see! it is \[\frac{8+\sqrt{x}}{\sqrt{8+x}}\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

yea hah i had no idea how to use the equation tool on computer

OpenStudy (anonymous):

if so you get \[\frac{8+\sqrt{1}}{\sqrt{8+1}}=\frac{8+1}{\sqrt{9}}\]

OpenStudy (anonymous):

\[=\frac{9}{3}=3\]

OpenStudy (anonymous):

i am practicing my latex

OpenStudy (anonymous):

yup..im happy i got the answer correcly

OpenStudy (anonymous):

write symbols that way

OpenStudy (anonymous):

ok will do